OpenStudy (anonymous):

. Choose the value for x that makes the open sentence true. 5 + 4 ● x = 32 + x2 (Points : 1) 5 4 3 2

OpenStudy (anonymous):

move the numbers to one side

OpenStudy (anonymous):

x^2 - 4x + 27(32-5) = 0

OpenStudy (anonymous):

quadratic equation formula ax^2 + bx + c = 0

OpenStudy (anonymous):

x = (-b (+/-) sqrt(b^2 - 4ac))/2a (plus or minus)

OpenStudy (anonymous):

a = 1, b = -4, c = 27

OpenStudy (anonymous):

x = (4 + sqrt(16 - 108)) / 2 x = (4 - sqrt(16 - 108)) / 2 16 - 108 = -92 square root of a negative number is "complex", signified by an "i" in the answer

OpenStudy (anonymous):

x = ( 4 + sqrt(92)i ) / 2 x = ( 4 - sqrt(92)i ) / 2

OpenStudy (anonymous):

x = ( 4 - 9.59i ) / 2 x = ( 4 + 9.59i ) / 2

OpenStudy (anonymous):

x = ( -5.59i ) / 2 x = ( 13.59i ) / 2

OpenStudy (anonymous):

x = -2.79i or 6.795 (you always have 2 answers for quadratic equations)

OpenStudy (anonymous):

2

OpenStudy (anonymous):

4 my bad

OpenStudy (anonymous):

so its 4.... did i help?

OpenStudy (anonymous):

good :)

OpenStudy (anonymous):

:] THANK YOU

OpenStudy (anonymous):

np